OpenStudy (anonymous):

Medal Math help Solve for x: −3|x + 7| = −12 x = 5/3, x = −19/3 x = −3, x = −11 x = −3, x = 11 No solution

OpenStudy (lynfran):

first step divide both sides by -3...what do u get?

OpenStudy (anonymous):

so divide -3 to -3

OpenStudy (lynfran):

yes both sides of the equation

OpenStudy (anonymous):

1

OpenStudy (domebotnos):

Did you deduct 7 from both sides? Remember PEMDAS. First in the parentheses.

OpenStudy (lynfran):

yes the -3 will cancel out on the left side of the equation...now what do u have remaining

OpenStudy (anonymous):

sorry not understanding.

OpenStudy (domebotnos):

You must do what's in the parentheses first. P () E x^2 M • D / A+ S -

OpenStudy (domebotnos):

Go by that list.

OpenStudy (domebotnos):

So you must deduct 7 from 7 and 13

OpenStudy (domebotnos):

-12, my bad

OpenStudy (anonymous):

x + 7

OpenStudy (lynfran):

-3|x+7|=-12 |x+7|=-12/-3 |x+7|=4 (x+7)=4 and (x+7)=-4

OpenStudy (lynfran):

solve for x

OpenStudy (anonymous):

when i was watching a video on this its says grab the -3 and put it under the 7 so is that right

OpenStudy (anonymous):



OpenStudy (lynfran):

x+7=4 x=4-7 x=-3 x+7=-4 x=-4-7 x=-11

OpenStudy (anonymous):

so -3 + 7=4

OpenStudy (lynfran):

yes

OpenStudy (anonymous):

so all i had to do was use  the hole time